OpenStudy (samigupta8):

Let A={A:sinA =tanA} and B={A:cosA=1} be two sets.Then: A is a subset of B ...How??

OpenStudy (samigupta8):

My ans is B is a subset of A

OpenStudy (michele_laino):

I got these solution sets: \[\large \begin{gathered} A = \left\{ {x \in \mathbb{R}|x = k\pi ,\quad k \in \mathbb{Z}} \right\} \cup \left\{ {x \in \mathbb{R}|x = 2h\pi ,\quad h \in \mathbb{Z}} \right\} - \hfill \\ - \left\{ {x \in \mathbb{R}|x = \frac{\pi }{2} + q\pi ,\quad q \in \mathbb{Z}} \right\} \hfill \\ \hfill \\ B = \left\{ {x \in \mathbb{R}|x = 2k\pi ,\quad k \in \mathbb{Z}} \right\} \hfill \\ \end{gathered} \]

OpenStudy (samigupta8):

That clearly means the same as i came up with ... Ur ans is also B is a subset of A

OpenStudy (michele_laino):

yes!

OpenStudy (samigupta8):

Ans key says that A is a subset of B

OpenStudy (michele_laino):

sincerely I don't know what to say

OpenStudy (samigupta8):

Ans key is wrong perhaps!

OpenStudy (michele_laino):

please note that: the solutions we have this: \[\begin{gathered} \sin \theta = \tan \theta \Rightarrow \hfill \\ \hfill \\ \Rightarrow \sin \theta \left( {1 - \frac{1}{{\cos \theta }}} \right) = 0 \Rightarrow \hfill \\ \hfill \\ \Rightarrow \sin \theta \left( {\frac{{\cos \theta - 1}}{{\cos \theta }}} \right) = 0 \Rightarrow \hfill \\ \hfill \\ \Rightarrow \sin \theta = 0 \cup \cos \theta - 1 = 0 \hfill \\ \end{gathered} \] for first equation

OpenStudy (michele_laino):

oops.. I meant ...the solutions we have are these...

OpenStudy (michele_laino):

whereas for second equation are these ones: \[\cos \theta - 1 = 0\]

OpenStudy (samigupta8):

Exactly i too solve in this pattern only

OpenStudy (samigupta8):

Solved *

OpenStudy (samigupta8):

That clearly means that A can be 0,π,2π,....nπ for set A Whereas for set B we just have A=0,2π,4π,....,2nπ

OpenStudy (michele_laino):

I remembered when I learned for first time the measure theory, so in the topological sense, we can say that set A is more fine than the set B

OpenStudy (michele_laino):

nevertheless I don't think that it can be the needed explanation

OpenStudy (samigupta8):

So it is clear than B has got to be the subset of A and not the reverse of it

OpenStudy (michele_laino):

in the set theory I think so!

OpenStudy (michele_laino):

please, ask to another helper too

OpenStudy (samigupta8):

Who??

OpenStudy (michele_laino):

I think that we can try to get help from @ganeshie8

OpenStudy (michele_laino):

also @mayankdevnani can help us, I think!

ganeshie8 (ganeshie8):

I think it is sufficient to prove below : \[\cos A = 1~~\implies~~ \sin A = \tan A \]

OpenStudy (samigupta8):

Yeah! That again means the same...

ganeshie8 (ganeshie8):

Is that hard to prove ?

OpenStudy (mayankdevnani):

but converse is also true !

ganeshie8 (ganeshie8):

converse isn't true

OpenStudy (samigupta8):

I didn't get you @ganeshie8 .. Are u agreeing with ans that A is a subset of B ?

OpenStudy (mayankdevnani):

Answer is correct !

OpenStudy (mayankdevnani):

A is a subset of B

OpenStudy (mayankdevnani):

i have proved it !

OpenStudy (samigupta8):

How can we say than A is a subset of B when it is very clear that A contains more no . of elements than that in B?

OpenStudy (mayankdevnani):

Wait,lemme post the solution

OpenStudy (mayankdevnani):

Set A :- Equate the general solutions:- \[\large \bf n \pi +(-1)^n \alpha=n \pi+\alpha\] \[\large \bf (-1)^n \alpha-\alpha=0\] \[\large \bf \alpha[(-1)^n-1]=0\] \[\large \bf \alpha=0~or~(-1)^n=1\] where, n=0,2,4,6...... Set B :- Equate the general solution :- \[\large \bf cosA=1=\cos0\] \[\large \bf cosA~where~A=2n \pi \pm \alpha\] \[\large \bf A=2n \pi ~and~\alpha=0\] where, n=0,1,2,3,4,..........

OpenStudy (mayankdevnani):

We can clearly see that A is a subset of B by noticing the `n`

OpenStudy (mayankdevnani):

@ganeshie8 @Michele_Laino i think it might helps !

OpenStudy (samigupta8):

This method works only when we are given like sinC =sinD then only can we write as C=nπ+(-1)^n (D) N furthermore why can't we do it as suggested by @michele_laino It was the appropriate method too...

OpenStudy (samigupta8):

If u do it in that manner u will also come with the same ans as he and i got

OpenStudy (mayankdevnani):

What !

OpenStudy (mayankdevnani):

we have given sinA=tanA

OpenStudy (mayankdevnani):

in LHS, i wrote the general solution of sin in RHS, i wrote the general solution of tan

OpenStudy (samigupta8):

Yep that doesn't mean we can equate general solutions of the two ....

OpenStudy (mayankdevnani):

we can !

OpenStudy (mayankdevnani):

can you give me a good reason why we can't ?

ganeshie8 (ganeshie8):

I don't get why you guys are trying to solve the given equations

OpenStudy (mayankdevnani):

@ganeshie8 am i wrong in my method ?

ganeshie8 (ganeshie8):

One sec... Wouldn't you agree that proving `A is a subset of B` is same as proving the below statement : \[\cos A = 1~~\implies~~ \sin A = \tan A\] ?

OpenStudy (anonymous):

It's the definition of subset.\[ A\subseteq B \equiv \forall x \quad x\in A\implies x \in B \]

OpenStudy (samigupta8):

No, i think this implies B is a subset of A ...

OpenStudy (anonymous):

The equation ends up being: \[ \sin A = \tan A \implies \cos A = 1 \]But this is false, particularly when \(\sin A=0\).

ganeshie8 (ganeshie8):

I feel skughtly dyslexic at the moment, but do you see we don't have to solve anything here ?

OpenStudy (samigupta8):

Sir bt we can say that since set A has solutions {0,π,2π,3π,4π,.....} And set B has solutions {0,2π,4π,....} Which implies clearly that set B has got to be the subset of A and not the reverse of it...

OpenStudy (anonymous):

Suppose \(A=\pi\). Then \(\sin A = 0 = \tan A\), so it is in \(A\); hoever \(\cos A = -1 \neq 1\) so it is not in \(B\).

OpenStudy (samigupta8):

Exactly @wio

hartnn (hartnn):

sin A = tan A A = n pi A: 0,pi, 2pi, 3pi, .... cos A =1 A =2n pi B: 0, 2pi, 4pi, ... B is the subset of A

OpenStudy (samigupta8):

Good god! You agreed with my ans @hartnn

OpenStudy (samigupta8):

So it's set now that the ans key is wrong ...bcoz the final ans is this only : B is a subset of A .... Thanks guys...

OpenStudy (mayankdevnani):

i bet your answer is wrong !

ganeshie8 (ganeshie8):

Looks good to me

hartnn (hartnn):

mayank, you're saying 'n' CANNOT be = 1 here : \(\large \bf n \pi +(-1)^n \alpha=n \pi+\alpha\) but when \(\alpha = 0\), 'n' can indeed be = 1. note : you ALSO have a "OR" in your answer. ---> when you equate the general solutions, you only take those which are common to both functions!

OpenStudy (mayankdevnani):

hmm

hartnn (hartnn):

from \(\sin \theta = 0 \cup \cos \theta - 1 = 0\) also, it is very clear that set A has solutions for \(\sin \theta = 0\) more than that of set B.

OpenStudy (mayankdevnani):

thank you !

OpenStudy (mayankdevnani):

well i applied that concept first and when i red the answer is opposite, i changed my mind lol

hartnn (hartnn):

it is a good habit to challenge stuff mayank! either you learn new things or you teach things that are new to others :) keep it up! :)